# Distributed Blind Source Separation based on FastICA

### v1

@@ I Introduction
Distributed algorithms have been proposed to solve the ICA / blind source separation problem [12, 13]. In [12], the network-wide ICA problem is solved heuristically with a (suboptimal) divide and conquer-type approach based on local neighborhoods. In [13], the ICA problem is solved in a truly distributed fashion, using the well-known alternative direction method of multipliers (ADMM). However, this algorithm iterates over sample batches, requiring multiple retransmissions of (updated versions of) the same batch of samples at each node, which leads to large communication costs, often sharing even more data than what was collected at the node. Furthermore, for each new incoming batch of samples, the ADMM iterations have to start from scratch, making ADMM-based algorithms unfit for (adaptive) spatial filtering on streaming data [14]. Additionally, oscillatory convergence behavior was observed in [13] when considering distributed settings without a fusion center. Another potential approach would be to first perform a distributed dimensionality reduction algorithm and perform a centralized ICA on the projected data. However, such an approach is not always applicable and might lead to suboptimal results. For example, using a distributed principal component analysis procedure [15], we would select the subspace of sources with the highest variance when reducing dimensions. In low SNR scenarios, this high-variance subspace could mainly capture (Gaussian) noise sources, whereas ICA aims to extract the sources that are least Gaussian. This argument has been validated in practice, e.g., in the context of electroencephalography signals [16].
In this paper, we propose the DistrICA algorithm, a distributed algorithm that directly solves the network-wide ICA problem without first projecting the sensor data in a low-dimensional subspace. By only transmitting linearly fused sensor signals between the nodes, the communication burden is significantly reduced, while still obtaining a subset of the sources from the centralized ICA problem.
@@ II Problem Setting
where $A\in\mathbb{R}^{M\times M}$ is the mixture matrix, and $\mathbf{s}(t)\in\mathbb{R}^{M}$ is the vector containing statistically independent source signals $s_{m}(t)$, $m\in\{1,\dots,M\}$. Our objective is to recover the sources $s_{m}$ from observations of $\mathbf{y}$ by applying a linear filter $X^{*}$ such that $X^{*T}\mathbf{y}(t)=\mathbf{r}(t)$, where the entries of $\mathbf{r}(t)$ are equal to those of $\mathbf{s}(t)$ up to a scaling and permutation ambiguity [2]. Note that this can also be a partial recovery, in the sense that we might be only interested in a subset of the sources $s_{m}$, i.e., $\mathbf{r}(t)\in\mathbb{R}^{Q}$, with $Q\leq M$. There exists various methods for this purpose, such as maximizing non-Gaussianity, maximum likelihood estimation, or minimizing the mutual information [2, 17]. In this paper, we focus on the FastICA algorithm [1], which consists of first applying a pre-whitening step on $\mathbf{y}$ to obtain a signal $\mathbf{z}$, followed by an orthogonal transformation that maximizes the “non-Gaussianity” of the demixed signals [^1] The core idea behind this is that a mixture of non-Gaussian sources is closer to a Gaussian distribution than any of the unmixed sources.. Formally, the whitening procedure can be written as
$\mathbf{z}(t)=ED^{-1/2}E^{T}\mathbf{y}(t),$ (2)
where $E$ and $D$ are obtained from the eigenvalue decomposition of the covariance matrix of $\mathbf{y}$, i.e., $R_{\mathbf{yy}}=\mathbb{E}[\mathbf{y}\mathbf{y}^{T}]=EDE^{T}$, such that $\mathbb{E}[\mathbf{z}\mathbf{z}^{T}]=I$. The ICA filters are then obtained by solving the following optimization problem
@@
The function $F$ is usually chosen as $F(x)=\text{log cosh}(x)$ or $F(x)=-\text{exp}(-x^{2}/2)$, making $\mathbb{E}[F(x)]$ a proxy for the negentropy of the random variable $x$ (also known as the “non-Gaussianity”) [2]. The $Q$ least Gaussian independent components are then found by applying a solution $W^{*}$ of (3) to $\mathbf{z}:W^{*T}\mathbf{z}$. Equivalently, we can express this result using the original data $\mathbf{y}$ as $X^{*T}\mathbf{y}$, where $X^{*}=ED^{-1/2}E^{T}W^{*}$ from (2). The steps of the FastICA algorithm are presented in Algorithm 1, where $F^{\prime}$ and $F^{\prime\prime}$ correspond to the first and second order derivative of $F$.
Let us now consider a sensor network with $K$ nodes, within the node set $\mathcal{K}=\{1,\dots,K\}$, where the topology of the network is given by a graph $\mathcal{G}$. Each node $k$ measures an $M_{k}-$channel signal $\mathbf{y}_{k}$, such that, defining
@@
Solving distributed problems where spatial information is required is possible with the Distributed Adaptive Signal Fusion (DASF) algorithm [14, 19], however the ICA problem does not trivially fit the family of problems that the DASF algorithm can solve. Nevertheless, in the next section, we will explain how the two-step problem (2)-(3) can be cast into the DASF framework, and in particular, how the across-node pre-whitening step (2) can be bypassed, such that only per-node pre-whitening operations are required.
@@
$\mathbf{z}(t)\leftarrow ED^{-1/2}E^{T}\mathbf{y}(t)$
@@ III The Distributed ICA algorithm
In this section, we present a method based on the DASF framework [14] to solve (2)-(3) in a distributed and adaptive fashion with provable convergence guarantees [19, 20].
@@ III-A Reformulating ICA as a DASF problem
$\underset{X\in\mathbb{R}^{M\times Q}}{\text{min. }}$ $\mathbb{E}[G(X^{T}\mathbf{y})]$ (5)
s.t. $\mathbb{E}[H_{j}(X^{T}\mathbf{y}(t))]\leq 0\;\textrm{ $\forall j\in\mathcal{J}_{I}$,}$
$\mathbb{E}[H_{j}(X^{T}\mathbf{y}(t))]=0\;\textrm{ $\forall j\in\mathcal{J}_{E}$,}$
where, $H_{j}$’s denote constraint functions of the problem and the sets $\mathcal{J}_{I}$ and $\mathcal{J}_{E}$ correspond to index sets of inequality and equality constraints respectively. Note that the family of the problems fitting the DASF framework is larger than the one represented by the formulation in (5), which is however omitted here for conciseness. Let us now cast Problem (2)-(3) into a problem fitting the DASF framework by embedding the pre-whitening step (2) within the ICA optimization problem (3), by making the change of variables: $X=ED^{-1/2}E^{T}W$, resulting in the equivalent problem
@@ III-B Data flow of DistrICA
In the distributed setting introduced in Section II, let every node $k\in\mathcal{K}$ have an estimate $X_{k}^{i}$ at iteration $i$ of the block $X_{k}\in\mathbb{R}^{M_{k}\times Q}$ of $X$, partitioned in the same way as $\mathbf{y}$ in (4):
@@
where $\mathcal{N}_{k}$ denotes the set of neighboring nodes of node $k$ after pruning, and $\widehat{\mathbf{y}}_{k\rightarrow n}^{i}\in\mathbb{R}^{Q}$ is the data transmitted by node $k$ to its neighboring node $n$. Note that the second term of (10) is recursive and vanishes for the leaf nodes of the tree, which initialize the process such that the inward flow naturally arises without central coordination (a node $l$ sends its data to node $n$ from the moment it has received data from all its neighbors except one, being node $n$). In this way, each node effectively transmits a $Q-$channel signal, independent of the size of the network, making the communication bandwidth fully scalable.
@@
from all its neighbors $n\in\mathcal{N}_{q}$, where $\mathcal{B}_{nq}$ is the branch of $\mathcal{T}^{i}(\mathcal{G},q)$ that is connected to node $q$ via node $n$ (excluding node $q$ itself).
@@ III-C Updating step
With the data exchange described in the previous subsection, node $q$ is now in possession of $N$ samples of $\widehat{\mathbf{y}}_{n\rightarrow q}^{i}$ for $n\in\mathcal{N}_{q}$ together with $N$ samples of its own (uncompressed) signal $\mathbf{y}_{q}$. Stacking these, we define the data available at node $q$ and iteration $i$ as follows:
@@
where $\widetilde{M}_{q}=|\mathcal{N}_{q}|\cdot Q+M_{q}$. At this step, the DASF algorithm [14] requires the updating node $q$ to use the signal available at node $q$, i.e., $\widetilde{\mathbf{y}}_{q}^{i}$ to solve the following compressed version of the original problem (6):
@@
where $R^{i}_{\widetilde{\mathbf{y}}_{q}\widetilde{\mathbf{y}}_{q}}=\mathbb{E}[\widetilde{\mathbf{y}}_{q}^{i}\widetilde{\mathbf{y}}_{q}^{iT}]$ is the covariance matrix of $\widetilde{\mathbf{y}}_{q}^{i}$. As seen in (13), a new variable $\widetilde{X}_{q}$ is then defined to act in an analogous way to $X$ locally. Remarkably, (13) has the exact same form as (6), and can therefore be solved locally at node $q$ by applying the FastICA algorithm described in Algorithm 1, based on the batch of $N$ available samples of $\widetilde{\mathbf{y}}_{q}^{i}$. Note that we have thus effectively bypassed the network-wide pre-whitening step of (2) and replaced it with a local pre-whitening step instead at node $q$.
@@
      2) Every node $k$ collects $N$ samples of $\mathbf{y}_{k}$. All nodes compress these to $N$ samples of $\widehat{\mathbf{y}}^{i}_{k}$.
      3) The nodes sum-and-forward their compressed data towards node $q$ via the recursive rule (10). Node $q$ eventually receives $N$ samples of $\widehat{\mathbf{y}}^{i}_{n\rightarrow q}$ given in (11), from all its neighbors $n\in\mathcal{N}_{q}$.
@@
such that each partition has a corresponding block in the partitioning of $\widetilde{\mathbf{y}}_{q}^{i}$ as defined in (12). The $Q$ least Gaussian sources can then be extracted at node $q$ by computing $\widetilde{X}_{q}^{*T}\widetilde{\mathbf{y}}^{i}_{q}=X^{(i+1)T}\mathbf{y}$, and can be transmitted from node $q$ to other nodes acting as data sinks if necessary. Additionally, the blocks $G_{n}^{i+1}$ are disseminated into the corresponding subgraph $\mathcal{B}_{nq}$ through node $n$. This allows every node to update its local variable estimator as
@@
This procedure is then repeated with a new updating node (e.g., chosen in a round-robin fashion) and a new set of $N$ samples of $\mathbf{y}$ at each iteration, such that changes in the statistical properties can be tracked, making the algorithm adaptive. In particular, $X^{i}$ is an estimator of $X^{*}(t)$ for $t=iN$. The steps of the proposed DistrICA algorithm are summarized in Algorithm 2.
Since DistrICA is derived according to the technical principles of the DASF framework, its convergence is guaranteed by the theoretical results in [19]. Note that [19] imposes some mild – yet highly technical – conditions, which can be shown to hold for the case of the DistrICA, but which are omitted here for conciseness. One crucial condition is that the number of constraints in (5) has to be smaller than $Q^{2}$, which is satisfied since the number of constraints in (6) is equal to $Q(Q+1)/2$ due to the symmetry of $R_{\mathbf{yy}}$ in (6).
Note that since (6) is a non-convex problem, the results in [19] only guarantee convergence to a stationary point of (3), which is also the case for the centralized FastICA algorithm. Nevertheless, the FastICA algorithm almost always converges to a global optimum in practice [17], such that the same holds for the DistrICA algorithm.
@@ IV Simulations
We consider a wireless sensor network where each node has $M_{k}=5$ sensors, and therefore measure a $5-$channel local signal $\mathbf{y}_{k}$. Throughout this section, we consider networks randomly generated using the Erdős-Rényi model with connection probability equal to $0.8$ and take $F(x)=\text{log cosh}(x)$ in the objective of (6). The signal model is as given in (1) and generated as follows. The elements of the mixture matrix $A\in\mathbb{R}^{M\times M}$ are drawn independently at random from the standard Gaussian distribution, i.e., $\mathcal{N}(0,1)$. We consider $M$ independent sources in $\mathbf{s}$, where we take $s_{1}$ to be a sinusoid, and $s_{2}$ to be a square signal (i.e., $Q=2$), while $s_{m}$, $2<m\leq M$ are taken to be convex combinations of uniformly and normally distributed noise, i.e., $s_{m}(t)=\alpha_{m}u_{m}(t)+(1-\alpha_{m})n_{m}(t)$, where $u_{m}\sim\mathcal{U}[-0.5,0.5]$, $n_{m}\sim\mathcal{N}(0,1)$, and $\alpha_{m}\in[0,1]$ is different for each source $m$. The resulting mixtures observed at the different sensors are normalized to unit variance. The goal is to separate the two non-Gaussian target sources from the near-Gaussian (noise) sources. Note that all sources $\mathbf{s}$ have more or less the same variance, hence a principal component analysis will not be able to separate the target subspace from the noise subspace.
Figure 1: Normalized error for varying number of nodes $K$ (top) and sources $Q$ (bottom).
@@
To assess the convergence of the DistrICA algorithm to the centralized solution, we consider a stationary setting, where at each iteration, $N=10^{4}$ samples are measured at the sensor nodes, and the necessary statistical parameters of the stochastic signals are approximated using a temporal averaging over these samples. We use the normalized squared error to measure the accuracy of the estimator $\{X^{i}\}_{i}$ obtained through the DistrICA algorithm over iterations $i$, given by $\varepsilon(i)=\text{median}\left(\frac{\|X^{i}-X^{*}\|^{2}}{\|X^{*}\|^{2}}\right)$, where the median is taken over $30$ Monte-Carlo runs under the same experimental settings. An optimal solution $X^{*}$ of the centralized ICA problem (6) is used for comparison and obtained through the FastICA algorithm in a centralized setting, where the stopping criterion is chosen as a normed difference of $10^{-8}$ between two consecutive iterates of the variable $X$ and a maximum number of iterations equal to $1000$. The same stopping criterion is initially used for solving the compressed ICA problem (13) at each iteration of the DistrICA algorithm. All simulations have been performed using the DASF toolbox [21].
Figure 1 (top) shows the convergence plots for two different network sizes: $K=5$ and $K=8$. We observe that although both settings lead to convergence towards an optimal solution of the ICA problem, the setting with fewer number of nodes does so faster. This is expected, as the per-node updating frequency is lower in a larger network due to the sequential updating procedure. Furthermore, since $M_{k}=5$ in both networks, the larger network solves an ICA problem with larger dimensions. On the other hand, in Figure 1 (bottom), we fix $K=8$ and consider two different numbers of sources $Q$. In the case $Q=1$, the source of interest is a square signal while all other signals are generated using $s_{m}(t)=\alpha_{m}u_{m}(t)+(1-\alpha_{m})n_{m}(t)$ as in the previous experiment. Note that the case $Q=1$ implies that the compressed signals $\widehat{y}_{k}$ transmitted within the network are comprised of only a single channel. This explains the slower convergence obtained for $Q=1$ compared to $Q=2$ when the number of iteration grows, as it implies that the compression ratio $M_{k}/Q$ increases, therefore requiring a smaller communication burden within the network at the expense of a slower convergence.
@@
Figure 2 presents results of the DistrICA algorithm for the scenarios presented previously when the FastICA solver of the compressed problem (13) is not exact, i.e., where we only let the nodes partially solve their local compressed ICA problem. In this experiment, we stop the local FastICA algorithm when it reaches an error of $10^{-3}$ in the normed difference between two consecutive filter values or a maximum of $10$ iterations. This allows to reduce the computational burden at the nodes while still guaranteeing convergence [20]. We see that until a certain error level, the convergence is not affected by this partial updating scheme and follows similar convergence properties while greatly reducing the computations performed at each node.
Figure 2: Comparison of the error for DistrICA with full and partial solutions of the local compressed problems.

### v2

@@ I Introduction
Distributed algorithms have been proposed to solve the ICA / blind source separation problem [12, 13]. In [12], the network-wide ICA problem is solved heuristically with a (suboptimal) divide and conquer-type approach based on local neighborhoods. In [13], the ICA problem is solved in a truly distributed fashion, using the well-known alternative direction method of multipliers (ADMM). However, this algorithm iterates over sample batches, requiring multiple retransmissions of (updated versions of) the same batch of samples at each node, which leads to a large communication cost scaling poorly with the network size, often sharing even more data than what was collected at the node. Furthermore, for each new incoming batch of samples, the ADMM iterations have to start from scratch, making ADMM-based algorithms unfit for (adaptive) spatial filtering on streaming data [14]. We note that most distributed estimation methods based on ADMM, consensus, or diffusion [15, 16, 17, 18] were originally designed for a setting where the data is partitioned in the sample dimension, making them unfit for the distributed ICA problem, where the data is partitioned in the channel dimension. Although some of these methods can be adapted to a channel-partitioning setting, they would face similar challenges as the ADMM-based ICA method in [13], requiring multiple iterations over every new sample batch. Another potential approach would be to first perform a distributed dimensionality reduction algorithm and perform a centralized ICA on the projected data. However, this is not always applicable and might lead to suboptimal results. For example, using a distributed principal component analysis procedure [19], we would select the subspace of sources with the highest variance when reducing dimensions. In low SNR scenarios, this high-variance subspace could mainly capture noise sources, instead of the relevant underlying sources [20].
In this paper, we propose DistrICA, a distributed algorithm that directly solves the network-wide ICA problem in an adaptive setting with streaming data, without first projecting the sensor data in a low-dimensional subspace. By only sharing linearly fused sensor signals between the nodes, the communication burden is significantly reduced, while still obtaining a subset of the sources from the centralized ICA problem.
@@ II Problem Setting
where $A\in\mathbb{R}^{M\times M}$ is the mixture matrix, and $\mathbf{s}(t)\in\mathbb{R}^{M}$ is the vector containing statistically independent source signals $s_{m}(t)$, $m\in\{1,\dots,M\}$. In the remaining parts of this paper, we will omit the time index $t$ unless we want to explicitly refer to a specific time sample. Our objective is to recover the sources $s_{m}$ from observations of $\mathbf{y}$ by applying a linear filter $X^{*}$ such that $X^{*T}\mathbf{y}=\mathbf{r}$, where the entries of $\mathbf{r}$ are equal to those of $\mathbf{s}$ up to a scaling and permutation ambiguity [2]. Note that this can also be a partial recovery, in the sense that we might be only interested in a subset of the sources $s_{m}$, i.e., $\mathbf{r}\in\mathbb{R}^{Q}$, with $Q\leq M$. There exists various methods for this purpose, such as maximizing non-Gaussianity, maximum likelihood estimation, or minimizing the mutual information [2, 21]. In this paper, we focus on the FastICA algorithm [1], which consists of first applying a pre-whitening step on $\mathbf{y}$ to obtain a signal $\mathbf{z}$, followed by an orthogonal transformation that maximizes the “non-Gaussianity” of the demixed signals [^1] The core idea behind this is that a mixture of non-Gaussian sources is closer to a Gaussian distribution than any of the unmixed sources.. Formally, the whitening procedure can be written as
$\mathbf{z}=ED^{-1/2}E^{T}\mathbf{y},$ (2)
such that $\mathbb{E}[\mathbf{z}\mathbf{z}^{T}]=I$, where $E$ and $D$ are obtained from the eigenvalue decomposition of the covariance matrix of $\mathbf{y}$, i.e., $R_{\mathbf{yy}}=\mathbb{E}[\mathbf{y}\mathbf{y}^{T}]=EDE^{T}$, assumed to be full rank [^2] This is to guarantee well-posedness [22, 23] of the problem, which can be briefly summarized as requiring that a small change in the inputs of the problem (the data) implies a small change in the output (the optimal solution).. The ICA filters are then obtained by solving the following problem
@@
$F$ is often chosen as $F(x)=\text{log cosh}(x)$ or $F(x)=-\text{exp}(-x^{2}/2)$, making $\mathbb{E}[F(x)]$ a proxy for the negentropy of the random variable $x$ (also known as the non-Gaussianity) [2]. The $Q$ least Gaussian independent components are then found by applying a solution $W^{*}$ of (3) to $\mathbf{z}:W^{*T}\mathbf{z}$. Equivalently, we can express this result using the original data $\mathbf{y}$ as $X^{*T}\mathbf{y}$, where $X^{*}=ED^{-1/2}E^{T}W^{*}$ from (2). The steps of FastICA are presented in Algorithm 1, where $F^{\prime}$ and $F^{\prime\prime}$ correspond to the first and second order derivative of $F$.
Let us now consider a sensor network represented by a graph $\mathcal{G}$ with $K$ nodes within the set $\mathcal{K}=\{1,\dots,K\}$. Each node $k$ measures an $M_{k}-$channel signal $\mathbf{y}_{k}$, such that, defining
@@
$\mathbf{z}\leftarrow ED^{-1/2}E^{T}\mathbf{y}$
@@ III The Distributed ICA algorithm
In this section, we present a distributed ICA method based on the Distributed Adaptive Signal Fusion (DASF) framework [14, 25, 26, 27, 28], which allows solving (adaptive) channel-partitioned distributed optimization problems. DASF uses a data-driven fuse-and-forward approach, where the $N$ most recent samples of all nodes are forwarded towards an updating node, while linearly fusing them along the way. The updating node collects these fused streams and locally solves a compressed version of the original network-wide problem to find the in-network fusion rules for the next iteration. By rotating the updating node at each iteration, convergence to the optimal solution can be achieved [25, 26]. However, the ICA problem does not trivially fit the family of problems that the DASF algorithm can solve. Nevertheless, in Section III-A, we will explain how the two-step problem (2)-(3) can be cast into the DASF framework, and in particular, how the across-node pre-whitening step (2) can be bypassed, such that only per-node pre-whitening operations are required.
@@ III-A Reformulating ICA as a DASF problem
[…]
$\mathbb{E}[H_{j}(X^{T}\mathbf{y})]=0\;\textrm{ $\forall j\in\mathcal{J}_{E}$,}$
where the $H_{j}$’s denote constraint functions of the problem and the sets $\mathcal{J}_{I}$ and $\mathcal{J}_{E}$ correspond to index sets of inequality and equality constraints respectively. The DASF algorithm can solve problems of the form (5) in a fully distributed and time-adaptive fashion with provable convergence guarantees [25]. Note that the family of the problems fitting the DASF framework is larger than the one represented by the formulation in (5), which is omitted here for conciseness. Let us now cast (2)-(3) into a problem fitting the DASF framework by embedding the pre-whitening step (2) within the ICA optimization problem (3), by making the change of variables: $X=ED^{-1/2}E^{T}W$, resulting in the equivalent problem
@@ III-B Data flow of DistrICA
Within the distributed problem setting introduced in Section II, let every node $k\in\mathcal{K}$ have an estimate $X_{k}^{i}$ at iteration $i$ of the block $X_{k}\in\mathbb{R}^{M_{k}\times Q}$ of $X$, partitioned as $\mathbf{y}$ in (4):
@@
where $\mathcal{N}_{k}$ denotes the set of neighboring nodes of node $k$ after pruning, and $\widehat{\mathbf{y}}_{k\rightarrow n}^{i}\in\mathbb{R}^{Q}$ is the data transmitted by node $k$ to its neighboring node $n$. Note that the second term of (10) is recursive and vanishes for the leaf nodes of the tree, which initializes the process such that the inward flow naturally arises without central coordination (a node $l$ sends its data to node $n$ from the moment it has received data from all its neighbors except one, being node $n$). In this way, each node effectively transmits a $Q-$channel signal, independent of the size of the network, making the communication bandwidth fully scalable.
@@
from all its neighbors $n\in\mathcal{N}_{q}$, where $\mathcal{B}_{nq}$ is the branch of $\mathcal{T}^{i}(\mathcal{G},q)$ that is connected to $q$ via $n$ (excluding $q$ itself).
@@ III-C Updating step
With the data exchange described in the previous subsection, node $q$ is now in possession of $N$ samples of $\widehat{\mathbf{y}}_{n\rightarrow q}^{i}$ for $n\in\mathcal{N}_{q}$ and $N$ samples of its own (uncompressed) signal $\mathbf{y}_{q}$. Stacking these, we define the data available at node $q$ and iteration $i$ as
@@
where $\widetilde{M}_{q}=|\mathcal{N}_{q}|\cdot Q+M_{q}$. At this step, the DASF algorithm [14] requires the updating node $q$ to use the available signal, i.e., $\widetilde{\mathbf{y}}_{q}^{i}$ to solve the following compressed version of (6):
@@
where $R^{i}_{\widetilde{\mathbf{y}}_{q}\widetilde{\mathbf{y}}_{q}}=\mathbb{E}[\widetilde{\mathbf{y}}_{q}^{i}\widetilde{\mathbf{y}}_{q}^{iT}]$ is the covariance matrix of $\widetilde{\mathbf{y}}_{q}^{i}$. As seen in (13), a new variable $\widetilde{X}_{q}$ is defined to act analogously to $X$ locally. Remarkably, (13) has the exact same form as (6), and can therefore be solved locally at node $q$ by applying the FastICA algorithm described in Algorithm 1, based on the batch of $N$ available samples of $\widetilde{\mathbf{y}}_{q}^{i}$. We have thus effectively bypassed the network-wide pre-whitening step (2) and replaced it with a local one instead at node $q$.
@@
      2) All nodes $k$ collect $N$ samples of $\mathbf{y}_{k}$ and compress them into $\widehat{\mathbf{y}}^{i}_{k}$.
      3) The nodes sum-and-forward $\widehat{\mathbf{y}}^{i}_{k}$ towards node $q$ via the recursive rule (10). Node $q$ eventually receives $N$ samples of $\widehat{\mathbf{y}}^{i}_{n\rightarrow q}$ given in (11), from all its neighbors $n\in\mathcal{N}_{q}$.
@@
such that each partition has a corresponding block in the partitioning of $\widetilde{\mathbf{y}}_{q}^{i}$ as defined in (12). The $Q$ least Gaussian sources can then be extracted at node $q$ by computing $\widetilde{X}_{q}^{*T}\widetilde{\mathbf{y}}^{i}_{q}=X^{(i+1)T}\mathbf{y}$, and be transmitted from node $q$ to other nodes acting as data sinks if necessary. The blocks $G_{n}^{i+1}$ are disseminated into the corresponding subgraph $\mathcal{B}_{nq}$ through node $n$, allowing every node to update its local estimator as
@@
This procedure is then repeated with a new updating node (e.g., chosen in a round-robin fashion) and a new set of $N$ samples of $\mathbf{y}$ at each iteration, such that changes in the statistical properties can be tracked, making the algorithm adaptive. In particular, $X^{i}$ is an estimator of $X^{*}(t)$ for $t=iN$. The steps of the proposed DistrICA algorithm are summarized in Algorithm 2. At the updating node, the computational complexity of DistrICA is in the order of the one of FastICA applied on the compressed data $\widetilde{\mathbf{y}}_{q}$, namely $\mathcal{O}(\widetilde{M}_{q}^{2}(N+\widetilde{M}_{q})+QT(\widetilde{M}_{q}N+\widetilde{M}_{q}^{2}Q))$, where $T$ corresponds to the number of iterations of the FastICA loop.
Since DistrICA is derived according to the technical principles of the DASF framework, its convergence is guaranteed by the theoretical results in [25]. Note that [25] imposes some mild – yet highly technical – conditions, which can be shown to hold for the case of DistrICA, but which are omitted here for conciseness. One crucial condition is that the number of constraints in (5) has to be smaller than $Q^{2}$, which is satisfied since there are $Q(Q+1)/2$ constraints in (6) due to the symmetry of $R_{\mathbf{yy}}$. Additionally, we exploit the fact that the solution set of FastICA is finite, where solutions can only differ up to a sign change of their columns, instead of an invariance to scaling and permutations, guaranteeing convergence to a single point [25, Theorem 5]. As (6) is a non-convex problem, the results in [25] only guarantee convergence to a stationary point of (3), which is also the case for the centralized FastICA algorithm. Nevertheless, FastICA almost always converges to a global optimum in practice [21], such that the same holds for DistrICA.
Figure 1: (Top) Normalized error for varying number of nodes $K$. (Middle) Convergence in an adaptive setting. (Bottom) Comparison of the error for DistrICA with full and partial solutions.
@@ IV Simulations
We consider a wireless sensor network where each node has $M_{k}=5$ sensors, and therefore measure a $5-$channel local signal $\mathbf{y}_{k}$. Throughout this section, we consider networks randomly generated using the Erdős-Rényi model with connection probability equal to $0.8$ and take $F(x)=\text{log cosh}(x)$ in the objective of (6). The signal model is as given in (1) and generated as follows. The elements of the mixture matrix $A\in\mathbb{R}^{M\times M}$ are drawn independently at random from the standard Gaussian distribution, i.e., $\mathcal{N}(0,1)$. We consider $M$ independent sources in $\mathbf{s}$, where $s_{1}$ is a sinusoid, and $s_{2}$ a square signal (i.e., $Q=2$), while $s_{m}$, $2<m\leq M$ are convex combinations of uniformly and normally distributed noise, i.e., $s_{m}(t)=\alpha_{m}u_{m}(t)+(1-\alpha_{m})n_{m}(t)$, where $u_{m}\sim\mathcal{U}[-0.5,0.5]$, $n_{m}\sim\mathcal{N}(0,1)$, and $\alpha_{m}\in[0,1]$ is different for each source $m$. The resulting mixtures observed at the different sensors are normalized to unit variance. The goal is to separate $s_{1}$ and $s_{2}$ from the near-Gaussian (noise) sources. Note that all sources $\mathbf{s}$ have more or less the same variance, hence a principal component analysis will not be able to separate the target subspace from the noise subspace.
@@
To assess the convergence of DistrICA to the centralized solution, we first consider a stationary setting, where at each iteration, $N=10^{4}$ samples are measured at the sensor nodes, and the necessary statistical parameters of the stochastic signals are approximated using a temporal averaging over these samples. We use the normalized squared error to measure the accuracy of the estimator $\{X^{i}\}_{i}$ obtained through DistrICA over iterations $i$, given by $\varepsilon(i)=\text{median}\left(\frac{\|X^{i}-X^{*}\|^{2}}{\|X^{*}\|^{2}}\right)$, where the median is taken over $30$ Monte-Carlo runs under the same experimental settings. An optimal solution $X^{*}$ of the centralized ICA problem (6) is used for comparison and obtained through FastICA in a centralized setting. All simulations have been performed using the DASF toolbox [29].
Fig. 1 (Top) shows convergence plots for two different network sizes: $K=5$ and $K=8$. We observe that, although both settings lead to convergence towards an optimal solution of ICA, the setting with fewer nodes does so faster. This is expected, as the per-node updating frequency is lower in a larger network due to the sequential updating procedure.
In Fig. 1 (Middle), we show the convergence of DistrICA in an adaptive setting, where the mixture matrix $A$ changes in time. For each new incoming batch of samples, $A$ is updated to $A+\Delta\cdot p(i)$, where $p$ is a scalar function defined by the red curve in Fig. 1, changing at each iteration $i$, i.e., every batch of $N$ samples will have a different mixture matrix. The entries of $\Delta$ are drawn from $\mathcal{N}(0,1)$ first, and then scaled such that $\|\Delta\|_{F}=0.005\cdot\|A\|_{F}$. To make the DistrICA algorithm more adaptive, we now also implement the alternative strategy from [30], which reuses sample batches $R$ times over multiple iterations (see [30] for details, $R=1$ corresponds to the original algorithm). Note that, in contrast to the stationary setting, we now observe a tracking error that saturates at a non-zero value due to changing signal statistics.
@@
Fig. 1 (Bottom) presents results of DistrICA when the FastICA solver of the compressed problem (13) is not exact, i.e., where we only let the nodes partially solve their local compressed ICA problem. In this experiment, we stop the local FastICA algorithm when it reaches an error of $10^{-3}$ in the normed difference between two consecutive filter values or a maximum of $10$ iterations. This allows reducing the computational burden at the nodes while still guaranteeing convergence [26]. We see that until a certain error level, the convergence is not affected by this partial updating scheme and follows similar convergence properties while greatly reducing the number of computations performed at each node.